1. If given a piece of "pure gold" jewelry, explain how you might determine if the piece is made of pure gold or not.

2. Acetone, a common solvent, has a density of 0.79 g/mL at 20˚C.  What is the volume of 100.g of acetone at 20˚C?

3. If a sample has an experimental refractive index of 1.4285 at 25.0 C, what is its refractive index at 20.0 C?

4. As the density of a liquid medium increases, would you expect the refractive index to increase or decrease? Hint: Would you expect light to go faster or slower in a more dense liquid medium

  1. What would be the effect on the observed melting point if sample were:
    1. too small –
    2. too large –
    3. poorly packed –
    4. heated too rapidly –
Answer #1

1. An acid test can be used to identify the authenticity of the gold . After a small scratch being made on to the jewelry, Drop of nitric acid can put and if no reaction occurs it's pure gold. Pure gold does not react with nitric acid.

2. Density = mass/ volume

Mass= 100g , Density = 0.79g/mL

Volume = mass/ density

= 100/0.79 = 126.58 mL

3. Refractive index inversely proportional to temperature

That is R1/R2 = T2/ T1

R1 = 1.4285 , R2 = ? T1 = 25°C , T2 = 25°C

R2 = R1*T1/T2

= 1.4385*25/20 = 1.7981

4. Refractive index = speed of light/ speed of light in medium

In a dense medium the speed of light decreases, that is refractive index becomes higher

a. Too small - no change in melting point since it does not depend on sample size

b. Too large - no change

c. Poorly packed - decrease in melting point

d. Rapid heating - no change since melting point is independent of heating rate